OpenStudy (anonymous):

if f(x) = 7(x-2)+4(x+1) , what is f (2) ?

OpenStudy (anonymous):

Sub x = 2 So answer is 12

OpenStudy (anonymous):

thank youuu :)

OpenStudy (anonymous):

no problem (:

OpenStudy (anonymous):

f(2)=7(2-2)+4(2+1) =7(0)+4(3) =0+12 =12